OpenStudy (anonymous):

pleeeeeaaase help 5/x^2-4+2/2/x=2/x-2

OpenStudy (anonymous):

write the question properly pls .... using eq notation..

OpenStudy (anonymous):

\[\frac{-4 x^2+x+5}{x^2}=-\frac{2 (x-1)}{x} \]?

OpenStudy (anonymous):